(1+x) -2 =? Explain it. -----------------------------------------

(1+x)-2=?
Explain it.
-----------------------------------------

Grade:10

1 Answers

Arun
5 years ago

binomial expansion of(1+x)^n , where n is a negative power,

(1+x)^-2 = 1+nx +n(n-1)x^2/2! +n(n-1)(n-2)x^3/3! +… an infinite series true for |x| less than 1.

=1+(-2)x +(-2)(-3)x^2/2! +(-2)(-3)(-4)x^3/3! +…

=1–2x +3x^2 -4x^3 +…

Regards

Arun (askIITiams forum expert)